# OccLE: Label-Efficient 3D Semantic Occupancy Prediction

### v1

@@ 3.3.2 Semi-Supervised Pipeline
We classify image augmentations by their effects on features. Brightness, contrast, and HSV adjustments modify global characteristics and constitute weak augmentation. Weather simulation and motion blur degrade feature clarity and serve as middle augmentation. Cutout removes entire regions of the feature map and thus represents strong augmentation. For LiDAR scans, we vary augmentation strength by adjusting the dropout probability, the dropout number, and the sampling number.
@@ 3.4 Semantic-Geometric Feature Grid Fusion
In this section, we fuse semantic and geometric feature grids to predict 3D semantic occupancy. Previous works [32, 34] employs attention mechanisms to enable long-range interactions within the feature grid. However, with sequence length ${HWZ}$, 3D DA still incurs substantial computational overhead. Recent work [36] adapts Mamba [17] module for efficient long-range dependency learning in this task. Inspired by this approach and given the spatial alignment of our semantic and geometric feature grids, we propose a Dual Mamba that process semantic and geometric feature grids independently, exchanges subsets of their channels for fusion. Specifically, we lift the feature map ${{\bf{F}}_{I}^{s}}$ to a feature grid via Equ. 2 and combine it with ${{\bf{F}}_{P}^{s}}$ to yield the semantic feature grid ${{\bf{F}}_{f}^{s}}$, Dual mamba then takes ${{\bf{F}}_{f}^{s}}$ and ${{\bf{F}}_{f}^{g}}$ as input and outputs the fused feature grid ${{\bf{F}}_{3d}^{f}}$. The detailed structure of Dual Mamba is illustrated in Appendix A.
@@
We use a 3D header and 2D header to predict the 3D semantic occupancy ${{\bf{Y}}}$ and the projected segmentation ${{\bf{y}}\in{{\mathbb{R}}^{h\times w}}}$ from feature grid ${{\bf{F}}_{3d}^{f}}$ and projected feature map ${{\bf{F}}_{2d}^{f}}$.
3.5 Losses and Training Strategy
3.5.1 Losses
@@
where ${{\ell_{ce}}}$ denotes the weighted cross-entropy loss, with weights computed from class frequencies. ${{\ell_{occ}}}$ is a comprehensive occupancy loss, following prior work [4, 37, 38]. ${{\bf{\bar{G}}}}$ represents the ground-truth of geometry. We apply ${{\ell_{occ}}\left({{\bf{S}},{\bf{\bar{S}}}}\right)}$ only to voxels containing more than one point, supervising voxel-level semantic learning.
3.5.2 Training Strategy
Input: Anotated dataset ${{{\cal D}_{l}}}$, unanotated dataset ${{{\cal D}_{u}}}$, Semantic branch ${{\Theta_{sem}}}$, Geometric branch ${{\Theta_{geo}}}$, Semantic-geometric feature grid fusion module ${{\Theta_{fus}}}$
Output: Trained ${{\Theta_{sem}}}$,${{\Theta_{geo}}}$,${{\Theta_{fus}}}$
Phase I: Train the semantic branch
Train ${{\Theta_{sem}}}$ with dataset ${{{\cal D}_{l}}}$,${{{\cal D}_{u}}}$, loss ${{\ell_{sem}}}$, label ${\{{\bf{\bar{S}}},{\bf{\bar{s}}}\}}$
Phase II: Train the geometric branch
Train the teacher model ${{\Theta_{geo-T}}}$ with dataset ${{{\cal D}_{l}}}$, loss ${{\ell_{sem}}}$, label ${\{{\bf{\bar{G}}}\}}$
Predict ${\{{{\bf{\bar{G}}}_{T}}\}}$ for dataset ${{{\cal D}_{u}}}$ via ${{\Theta_{geo-T}}}$
Train the student model ${{\Theta_{geo-S}}}$ with dataset ${{{\cal D}_{l}}}$,${{{\cal D}_{u}}}$, loss ${{\ell_{sem}}}$, label ${\{{\bf{\bar{G}}}\}}$,${\{{{\bf{\bar{G}}}_{T}}\}}$
Phase III: Train the semantic-geometric feature grid fusion module
Freeze ${{\Theta_{geo-S}}}$; Train ${{\Theta_{fus}}}$ and ${{\Theta_{sem}}}$ with dataset ${{{\cal D}_{l}}}$,${{{\cal D}_{u}}}$, loss ${{\ell_{fus}}}$,${{\ell_{sem}}}$, label ${\{{\bf{\bar{Y}}},{\bf{\bar{s}}}\}}$
Algorithm 1 The training strategy of OccLE
4 Experiment
4.1 Dataset and Metric
4.1.1 Dataset
We evaluate OccLE on the SemanticKITTI [27], which includes 22 outdoor driving scenarios: 10 sequences for training, 1 for validation, and 11 for testing. The dataset annotates scenes of size ${51.2m\times 51.2m\times 6.4m}$ with a voxel size of ${0.2m\times 0.2m\times 0.2m}$ for 20 semantic classes (19 semantics + 1 free). To simulate a label-efficient learning, we use only 10% of the voxel annotations from the training set, obtained via interval sampling.
4.1.2 Metric
4.2 Implementation Details
[…]
4.3 Main Result
We present a quantitative comparison on the SemanticKITTI validation set in Table 2. OccLE obtains competitive performance, achieving 16.59 % mIoU and 39.96 % IoU. It surpasses all fully supervised camera-based methods (HASSC [40] with 40.66 % mIoU) and approaches the performance of fully supervised multi-modal methods (OccLoff [22] with 22.62 % mIoU). As illustrated in Figure 4, OccLE achieves strong performance in complex scenes, accurately reconstructing buildings and vegetation, and distinguishing small objects such as traffic signs and poles.
Table 2: Quantitative results on SemanticKITTI validation set. Bold and underline represent the best and second best results, respectively.
@@ 4.4 Ablation Studies
We conduct all ablation studies on the validation set of the SemanticKITTI dataset. We focus on three aspects: semantic branch, geometric branch, and overall pipeline.
4.4.1 The Semantic Branch
To evaluate the extractor and pseudo label in the semantic branch, for comparison, we adopt ESPNetv2 [43] and MobileNet [44] as ${{{\cal G}_{s2d}}}$, employ 1-layer and 2-layer versions of sparse UNet3D as ${{{\cal G}_{s3d}}}$, and train using unaligned pseudo labels. As depicted in Table 3, 2D extractors designed for edge device suffer a notable performance degradation, whereas our method achieves the highest mIoU(${{\bf{s}}}$) while retaining remaining lightweight. Increasing the depth of the sparse UNet3D architecture improves mIoU(${{\bf{S}}}$) without substantially increasing computational cost. Moreover, aligned pseudo labels yield a significant mIoU gain by providing supervision for unaligned classes.
@@ 4.4.2 The Geometric Branch
To evaluate the effects of model architecture and training strategy in the geometric branch, we design a baseline that employs Conv3D and 3D DA for feature extraction and interaction, and compare our semi-supervision against full supervision. As depicted in Table 4(a), our cross-plane image and LiDAR feature synergy only has a minor precision drop while significantly reducing module inference time from 138.62 ms to 26.08 ms. Our data augmentation strategy boosts student model performance as augmentation degree increases. Under semi-supervision, our student model reaches 53.60 IoU, closely approaching the fully supervised performance of 56.77 IoU.
Table 4: Ablation studies for (a) geometric branch and (b) overall pipeline. “weak,” “medium,” and “strong” denote the degrees of data augmentation, -S and -T represents the student and teacher models. ‘Sem.’, ‘Geo.’, and ‘Proj.’ denote the semantic branch, geometric branch, and supervising with scatter-accumulated projection, respectively.
(a) Geometric branch.
Sup. Weak Middle Strong Model IoU
Full Ours-T 56.77
10% Baseline 52.79
10% Ours-T 52.66
10% Ours-S 52.86
10% Ours-S 52.88
10% Ours-S 53.60
(b) Overall pipeline.
@@
4.4.3 Overall Pipeline
As depicted in Table 4(b), compared to the baseline, our method increases mIoU from 9.64 to 16.59 using only 10% voxel annotations, closely approaching the fully supervised performance of 20.32 mIoU. Furthermore, ablating any module in OccLE results in a precision drop, demonstrating the essential role of these three modules in label-efficient learning. Additional ablation studies are presented in the Appendix C.2.
@@ 5 Conclusion
In this paper, we propose OccLE, a label-efficient 3D semantic occupancy prediction that leverages limited voxel annotations while preserving high performance. OccLE incorporates three key components: (1) distilling 2D foundation model to predict aligned pseudo label for supervising 2D and 3D semantic learning, (2) cross-plane image and LiDAR feature synergy for efficient geometry learning under semi-supervision, and (3) semantic-geometric feature grid fusion with Dual Mamba and scatter-accumulated projection for supervising unannotated regions. Experimental results demonstrate that OccLE achieves an mIoU of 16.59% on the SemanticKITTI validation set using only 10% of the voxel annotations.
@@ Appendix B Aligned Label Generation
We employ dataset-specific and open-vocabulary segmentation 2D foundation models to generate aligned pseudo labels. The class mappings are listed in Table B.1. We utilize SAM2 [8] to produce trunk class labels and integrate them into the MSeg [28] outputs. This example illustrates our aligned pseudo label generation process, which can scale to any dataset with different class taxonomies.
Table B.1: The aligned label generation.
@@
Appendix C Supplementary Experimental Results
C.1 Supplementary Ablation Study

### v2

@@ 3.3.2 Semi-Supervised Pipeline
We classify image augmentations by their effects on features. Brightness, contrast, and HSV adjustments modify global characteristics and constitute weak augmentation. Weather simulation and motion blur degrade feature clarity and serve as middle augmentation. Cutout removes entire regions of the feature map and thus represents strong augmentation. For LiDAR, we vary augmentation strength by adjusting the dropout probability, the dropout number, and the sampling number.
@@ 3.4 Semantic-Geometric Feature Grid Fusion
In this section, we fuse semantic and geometric feature grids to predict 3D semantic occupancy. Previous works [41, 43] employs attention mechanisms to enable long-range interactions within the feature grid. However, with sequence length HWZ{HWZ}, 3D DA still incurs substantial computational overhead. Recent work [20] adapts Mamba [8] module for efficient long-range dependency learning in this task. Inspired by this approach and given the spatial alignment of our semantic and geometric feature grids, we propose a Dual Mamba that process semantic and geometric feature grids independently, exchanges subsets of their channels for fusion. Specifically, we lift the feature map ${\bf{F}}_{I}^{s}$} to a feature grid via Equ. 2 and combine it with ${\bf{F}}_{P}^{s}$} to yield the semantic feature grid ${\bf{F}}_{f}^{s}$}, Dual mamba then takes ${\bf{F}}_{f}^{s}$} and ${\bf{F}}_{f}^{g}$} as input and outputs the fused feature grid ${\bf{F}}_{3d}^{f}$}. The detailed structure of Dual Mamba is illustrated in Appendix A.
@@
We use a 3D and 2D header to predict the 3D semantic occupancy $\bf{Y}$}} and the projected segmentation ${\bf{y}}\in{{\mathbb{R}}^{h\times w}}$} from feature grid ${\bf{F}}_{3d}^{f}$} and projected feature map ${\bf{F}}_{2d}^{f}$}.
3.5 Losses
@@
where $\ell_{ce}$}} denotes the weighted cross-entropy loss, with weights computed from class frequencies. $\ell_{occ}$}} is a comprehensive occupancy loss, following prior work [22, 17]. $\bf{\bar{G}}$}} represents the ground-truth of geometry. We apply ${\ell_{occ}}\left({{\bf{S}},{\bf{\bar{S}}}}\right)$} only to voxels with more than one point to supervise voxel-level semantics.
Table 2: Quantitative results on SemanticKITTI validation set. Bold and underline represent the best and second best results, respectively. Inp. and Sup. indicate the input modality and the supervision type, respectively.
@@
Table 3: Quantitative results on Occ3D-nuScenes validation set. $\dagger$} indicates the use of camera-visible masks during training.
Method Inp. Sup. mIoU
other
barrier
bicycle
bus
car
cons. veh.
motorcycle
pedestrian
traffic cone
trailer
truck
drive. surf.
other flat
sidewalk
terrain
manmade
vegetation
MonoScene [4] C Full 6.06 1.75 7.23 4.26 4.26 4.93 9.38 3.98 3.98 3.90 4.45 7.17 14.91 6.32 7.92 7.43 1.01 7.43
OccFormer [53] C Full 21.93 5.95 21.93 5.85 37.83 17.87 40.44 42.43 7.36 23.88 21.81 20.98 22.38 30.70 55.35 28.36 30.70 18.0
BEVFormer [26] C Full 26.88 5.94 30.29 12.32 34.40 39.17 14.44 16.45 17.22 9.27 13.90 26.36 50.99 30.96 34.66 22.73 6.76 6.97
TPVFormer [13] C Full 27.83 7.22 38.90 13.67 40.78 45.90 17.23 19.99 18.85 14.30 26.69 34.17 55.65 35.47 37.55 30.70 19.40 16.78
[…]
SelfOcc [14] C Self 9.30 0.00 0.15 0.66 5.46 12.54 0.00 0.80 2.10 0.00 0.00 8.25 55.49 0.00 26.30 26.54 14.22 5.60
OccNeRF [49] C Self 9.53 0.00 0.83 0.82 5.13 12.49 3.50 0.23 3.10 1.84 0.52 3.90 52.62 0.00 20.81 24.75 18.45 13.19
DistillNeRF [39] C Self 8.93 0.03 1.35 2.08 10.21 10.09 2.56 1.98 5.54 4.62 1.43 7.90 43.02 0.00 16.86 15.02 14.06 15.06
GaussTR [18] C Self 11.70 - 2.09 5.22 14.07 20.43 5.70 7.08 5.12 3.93 0.92 13.36 39.44 - 15.68 22.89 21.17 21.87
RenderOcc [31] C Self 23.93 5.69 27.56 14.36 19.91 20.56 11.96 12.42 12.14 14.34 20.81 18.94 68.85 33.35 42.01 43.94 17.36 22.61
OccLE (Ours) C+L 10% 27.53 28.58 0.00 27.33 37.44 8.13 0.00 17.76 0.00 12.56 20.18 71.51 30.97 38.46 38.94 27.90 24.37 83.95
4 Experiment
4.1 Dataset and Metric
4.1.1 Dataset
We evaluate OccLE on the SemanticKITTI [1] and Occ3D-nuScenes [37]. SemanticKITTI annotates scenes of size $51.2m\times 51.2m\times 6.4m$} with 0.2m{0.2m} voxel size for 20 classes (19 semantics + 1 free). Occ3D-nuScenes annotate the $80.0m\times 80.0m\times 6.4m$} with 0.4m{0.4m} voxel size for 18 classes (17 semantics + 1 free). To simulate a label-efficient learning, we use only 10% of the voxel annotations in the training set, obtained via interval sampling.
4.1.2 Metric
4.2 Implementation Details
All experiments are conducted on 2 NVIDIA A6000 GPUs with a batch size of 1. The AdamW optimizer is used with an initial learning rate of $2{e^{-4}}$} and a weight decay of $e^{-4}$}}, All training durations are set to 40 epochs, while the teacher model of the geometric branch is set to 100 epochs. More details are presented in Appendix D.2.
4.3 Main Result
We present a quantitative comparison on the SemanticKITTI validation set and Occ3D-nuScenes validation set in Table 2 and Table 3, respectively. For the SemanticKITTI dataset, OccLE obtains competitive performance, achieving 16.59 % mIoU and 39.96 % IoU. It surpasses all fully supervised camera-based methods (HASSC [40] with 40.66 % mIoU) and approaches the performance of fully supervised multi-modal methods (OccLoff [52] with 22.62 % mIoU). For the Occ3D-nuScenes dataset, OccLE is trained using camera-visible masks for 10% of the training samples and a single-frame input. The experimental results show that OccLE achieves 27.53 mIoU, outperforming all self-supervised methods (RenderOcc [31] with 23.93% mIoU) and approaching the performance of fully supervised methods. As illustrated in Figure 4, OccLE achieves strong performance in complex scenes, accurately reconstructing buildings and vegetation, and distinguishing small objects such as traffic signs and poles. Additional comparisons on the SemanticKITTI hidden test set and SSCBenchKITTI-360 [23] validation set are presented in Appendix D.3.
@@ 4.4 Ablation Studies
We conduct all ablation studies on the validation set of the SemanticKITTI dataset. Additional ablation studies are presented in the Appendix D.4.
@@
4.4.1 The Semantic Branch
To evaluate the extractor and pseudo label in the semantic branch, for comparison, we adopt ESPNetv2 [28] and MobileNet [12] as ${\cal G}_{s2d}$}}, employ 1-layer and 2-layer versions of sparse UNet3D as ${\cal G}_{s3d}$}}, and train using unaligned pseudo labels. As depicted in Table 4, 2D extractors designed for edge device exhibit a clear drop in performance, whereas our method achieves the highest mIoU($\bf{s}$}}) while remaining lightweight. Deeper sparse UNet3D architectures improves mIoU($\bf{S}$}}) without substantially increasing computational cost. Aligned pseudo labels significantly improve mIoU by offering supervision for unaligned classes.
@@ 4.4.2 The Geometric Branch
To evaluate the effects of model architecture and training strategy in the geometric branch, we design a baseline that employs Conv3D and 3D DA for feature extraction and interaction, and compare our semi-supervision against full supervision. As depicted in Table 5, our cross-plane image and LiDAR feature synergy only has a minor precision drop while significantly reducing module inference time from 138.62 ms to 26.08 ms. Our data augmentation strategy boosts student model performance as augmentation degree increases. Under semi-supervision, our student model reaches 53.60 IoU.
Table 5: Ablation study of the geometric branch, where “weak,” “medium,” and “strong” denote the degrees of data augmentation, -S and -T represents the student and teacher models.
Sup. Weak Middle Strong Model Inf. Time IoU
Full Ours-T 26.08 56.77
10% Baseline 138.62 52.79
10% Ours-T 26.08 52.66
10% Ours-S 26.08 52.86
10% Ours-S 26.08 52.88
10% Ours-S 26.08 53.60
4.4.3 Overall Pipeline
We use only 10% voxel annotations for supervision as the baseline. As depicted in Table 6, our method increases mIoU from 9.64 to 16.59 by incorporating a decoupled design and 2D-aligned pseudo labels, closely approaching the fully supervised performance of 20.32 mIoU. Using only aligned pseudo labels significantly reduces precision, indicating pattern collapse when voxel annotations are missing. Furthermore, ablating any module in OccLE results in a precision drop, demonstrating the essential role of these three modules in label-efficient learning.
Table 6: Ablation study of overall pipeline, where ‘Sem.’, ‘Geo.’, and ‘Self.’ denote the semantic branch, geometric branch, and self-supervision in the semantic-geometric fusion module, respectively.
@@ 4.4.3 Overall Pipeline
4.4.4 Voxel Annotation Ratio
As shown in Table 7, OccLE achieves 18.40% mIoU with only 15% voxel annotations, approaching the fully supervised camera+LiDAR method OccLoff [52] (22.62%). With just 5% of training data, it still reaches 13.11% mIoU, showing strong performance under limited supervision.
Table 7: Ablation study of voxel annotation ratio.
Ratio 5% 10% 15%
IoU 32.88 40.60 42.10
mIoU 13.11 16.59 18.40
@@ 5 Conclusion
In this paper, we propose OccLE, a label-efficient 3D semantic occupancy prediction that leverages limited voxel annotations while preserving high performance. OccLE incorporates three key components: (1) distilling 2D foundation model to predict aligned pseudo label for supervising 2D and 3D semantic learning, (2) cross-plane image and LiDAR feature synergy for efficient geometry learning under semi-supervision, and (3) semantic-geometric feature grid fusion with Dual Mamba and scatter-accumulated projection for supervising unannotated regions. Experiments show that OccLE achieves competitive performance with only 10% of voxel annotations on the SemanticKITTI and Occ3D-nuScenes datasets.
@@ Appendix B Aligned Label Generation
We employ dataset-specific and open-vocabulary segmentation 2D foundation models to generate aligned pseudo labels. The class mappings are listed in Table B.1 and Table B.2. We utilize SAM2 [32] to produce trunk class labels and integrate them into the MSeg [19] outputs. This example illustrates our aligned pseudo label generation process, which can scale to any dataset with different class taxonomies.
Table B.1: The aligned label generation on the SemanticKITTI dataset.
@@
Table B.2: The aligned label generation on the Occ3D-nuScenes dataset.
Occ3D-nuScenes Image Segmentation Occ3D-nuScenes Image Segmentation
Label Class Label Class Label Class Label Class
0 other rest sky et.al 8 traffic cone SAM2 traffic cone
1 barrier 130 streetlight 9 trailer 183 trailer
131 road barrier 10 truck 182 trailer
144 fence 11 drive. surf. 98 road
145 railing banister 12 other flat 96 playing field
145 guard rail 97 railroad
2 bicycle 175 bicycle 13 sidewalk 100 sidewalk pavemen
3 bus 180 bus 14 terrain 102 terrain
4 car 176 car 15 manmade 191 wall
5 cons. veh SAM2 cons. veh 192 window
6 motorcycle 178 motorcycle 132-141 traffic sign et al.
128 motorcyclist 16 vegetation 174 vegetation
7 pedestrian 125 person
Appendix C Training Strategy
Algorithm 1 The training strategy of OccLE
0:  Annotated dataset $\mathcal{D}_{l}$, unannotated dataset $\mathcal{D}_{u}$, semantic branch $\Theta_{\text{sem}}$, geometric branch $\Theta_{\text{geo}}$, fusion module $\Theta_{\text{fus}}$
0:  Trained $\Theta_{\text{sem}},\Theta_{\text{geo}},\Theta_{\text{fus}}$
1:  Phase I: Train the semantic branch Train $\Theta_{\text{sem}}$ with $\mathcal{D}_{l}$, $\mathcal{D}_{u}$, loss $\ell_{\text{sem}}$, label $\{\bar{\mathbf{S}},\bar{\mathbf{s}}\}$
2:  Phase II: Train the geometric branch Train teacher model $\Theta_{\text{geo-T}}$ with $\mathcal{D}_{l}$, loss $\ell_{\text{sem}}$, label $\{\bar{\mathbf{G}}\}$ Predict $\{\bar{\mathbf{G}}_{T}\}$ for $\mathcal{D}_{u}$ via $\Theta_{\text{geo-T}}$ Train student model $\Theta_{\text{geo-S}}$ with $\mathcal{D}_{l}$, $\mathcal{D}_{u}$, loss $\ell_{\text{sem}}$, labels $\{\bar{\mathbf{G}}\}$, $\{\bar{\mathbf{G}}_{T}\}$
3:  Phase III: Train the fusion module Freeze $\Theta_{\text{geo-S}}$ Train $\Theta_{\text{fus}}$ and $\Theta_{\text{sem}}$ with $\mathcal{D}_{l}$, $\mathcal{D}_{u}$, losses $\ell_{\text{fus}}$, $\ell_{\text{sem}}$, labels $\{\bar{\mathbf{Y}},\bar{\mathbf{s}}\}$
Appendix D Supplementary Experiments
Table D.3: Quantitative results on SemanticKITTI hidden test set. Bold and underline represent the best and second best results, respectively. Inp. and Sup. indicate the input modality and the supervision type, respectively.
Method Inp. Sup. IoU mIoU
car (3.92%)
bicycle (0.03%)
motorcycle (0.03%)
truck (0.16%)
other-veh. (0.20%)
person (0.07%)
bicyclist (0.07%)
motorcyclist (0.05%)
road (15.30%)
parking (1.12%)
sidewalk (11.13%)
other-grnd.(0.56%)
building (14.10%)
fence (3.90%)
vegetation (39.3%)
trunk (0.51%)
terrain (9.17%)
pole (0.29%)
traf.-sign (0.08%)
TPVFormer [13] C Full 34.25 11.26 19.20 1.00 0.50 3.70 2.30 1.10 2.40 0.30 55.10 27.40 27.20 6.50 14.80 11.00 13.90 2.60 20.40 2.90 1.50
OccFormer [53] C Full 34.53 12.32 21.30 1.50 1.70 3.90 3.20 2.20 1.10 0.20 55.90 31.50 30.30 6.50 15.70 11.90 16.80 3.90 21.30 3.80 3.70
VoxFormer [22] C Full 43.21 13.41 21.70 1.90 1.60 3.60 4.10 1.60 1.10 0.00 54.10 25.10 26.90 7.30 23.50 13.10 24.40 8.10 24.20 6.60 5.70
SurroundOcc [42] C Full 34.72 11.86 20.30 1.60 1.20 1.40 4.40 1.40 2.00 0.10 56.90 30.20 28.30 6.80 15.20 11.30 14.90 3.40 19.30 3.90 2.40
H2GFormer [41] C Full 43.52 14.60 23.70 0.60 1.20 5.20 5.00 1.10 0.10 0.00 57.90 30.00 30.40 6.90 24.00 14.60 25.20 10.70 25.80 7.50 7.10
SGN [29] C Full 45.42 15.76 25.40 4.50 0.90 4.50 3.70 0.50 0.30 0.10 60.40 28.90 31.40 8.70 28.40 18.10 27.40 12.60 28.40 10.00 8.30
LowRankOcc [54] C Full 38.47 13.56 20.90 3.30 2.70 2.90 4.40 2.40 1.70 2.30 52.80 25.10 27.20 8.80 22.10 14.40 22.90 8.90 20.80 7.00 7.00
Symphonies [17] C Full 42.19 15.04 23.60 3.60 2.60 3.20 5.60 3.20 1.90 2.00 58.40 26.90 29.30 11.70 24.70 16.10 24.20 10.00 23.10 7.70 8.00
HASSC [40] C Full 42.87 14.38 23.00 1.90 1.50 2.90 4.90 1.40 3.00 0.00 55.30 25.90 29.60 11.30 23.10 14.30 24.80 9.80 26.50 7.00 7.10
Bi-SSC [43] C Full 45.10 16.73 25.00 1.80 2.90 6.80 6.80 1.70 3.30 1.00 63.40 31.70 33.30 11.20 26.60 19.40 26.10 10.50 28.90 9.30 8.40
OccLE (Ours) C+L 10% 31.42 16.30 29.60 3.50 2.70 0.00 6.30 1.90 0.00 0.00 56.90 22.60 30.80 2.30 24.90 21.90 30.00 19.10 27.90 15.10 14.00
D.1 Supplementary Datasets
We evaluate OccLE on additional datasets: SSCBenchKITTI-360[23]. SSCBenchKITTI-360 shares its scene and voxel configuration with SemanticKITTI and annotates voxels for 19 class labels (18 Semantic + 1 Free).
D.2 Supplementary Implementation Details
In the semantic branch, the downsampling scale of ${\cal G}_{s2d}$}} and ${\cal G}_{s3d}$}} is $8\times$}; In the geometric branch, we utilize ResNet50 [10] and a 2-layer sparse Conv3D to extract features from image and LiDAR scans. In the semantic-geometric feature grid fusion module, we stack 4 Dual Mamba blocks. For the SemanticKITTI dataset, we adopt the multi-frame setting as in [22, 29, 41, 40], and crop the camera images to $1220\times 370$}. For the Occ3D-nuScenes dataset, we use a single-frame setting and crop the images to $1600\times 900$}. For the SSCBenchKITTI-360 dataset, we follow the same multi-frame setting as SemanticKITTI and crop the images to $1408\times 376$}.
D.3 Supplementary Quantitative Comparison
The quantitative comparison results on the SemanticKITTI hidden test set and SSCBenchKITTI-360 validation set are presented in Table D.3 and Table D.4. OccLE achieves 16.30 mIoU and 31.42 IoU on the SemanticKITTI hidden test set, ranking second among camera-based fully supervised methods. On the SSCBenchKITTI-360 validation set, OccLE achieves 16.38 mIoU and 52.44 IoU, performing competitively against both camera-based and LiDAR-based methods. These results indicate that OccLE can perform well even with limited voxel-level annotations.
Table D.4: Quantitative results on SSCBench-KITTI-360 validation set.
Method Inp. Sup. IoU mIoU
car (2.85%)
bicycle (0.01%)
motorcycle (0.01%)
truck (0.16%)
other-veh. (5.75%)
person (0.02%)
road (14.98%)
parking (2.31%)
sidewalk (6.43%)
other-grnd.(2.05%)
building (15.67%)
fence (0.96%)
vegetation (41.99%)
terrain (7.10%)
pole (0.22%)
traf.-sign (0.06%)
other-struct. (4.33%)
other-obj. (0.28%)
TPVFormer [13] C Full 40.22 13.64 21.56 1.09 1.37 8.06 2.57 2.38 52.99 11.99 31.07 3.78 34.83 4.80 30.08 17.52 7.46 5.86 5.48 2.70
OccFormer [53] C Full 40.27 13.81 22.58 0.66 0.26 9.69 3.82 2.77 54.30 13.44 31.53 3.55 36.42 4.80 31.00 19.51 7.77 8.51 6.95 4.60
VoxFormer [22] C Full 38.76 11.91 17.84 1.16 0.89 4.56 2.06 1.63 47.01 9.67 27.21 2.89 31.18 4.97 28.99 14.69 6.51 6.92 3.79 2.43
SGN [29] C Full 47.06 18.25 29.03 3.43 2.90 10.89 5.20 2.99 58.14 15.04 36.40 4.43 42.02 7.72 38.17 23.22 16.73 16.38 9.93 5.86
Symphonies [17] C Full 44.12 18.58 30.02 1.85 5.90 25.07 12.06 8.20 54.94 13.83 32.76 6.93 35.11 8.58 38.33 11.52 14.01 9.57 14.44 11.28
GaussianFormer [16] C Full 35.38 12.92 18.93 1.02 4.62 18.07 7.59 3.36 45.47 10.89 25.03 5.32 28.44 5.68 29.54 8.62 2.99 2.32 9.51 5.14
GaussianForme-2 [15] C Full 38.37 13.90 21.08 2.55 4.21 12.41 5.73 1.59 54.12 11.04 32.31 3.34 32.01 4.98 28.94 17.33 3.57 5.48 5.88 3.54
LMSCNet[33] L Full 47.53 13.65 20.91 0.00 0.00 0.26 0.00 0.00 62.95 13.51 33.51 0.20 43.67 0.33 40.01 26.80 0.00 0.00 3.63 0.00
SSCNet[36] L Full 53.58 16.95 31.95 0.00 0.17 10.29 0.58 0.07 65.70 17.33 41.24 3.22 44.41 6.77 43.72 28.87 0.78 0.75 8.60 0.67
OccLE (Ours) C+L 10% 52.44 16.38 31.14 0.34 0.28 4.66 1.39 2.05 53.47 9.15 29.65 4.32 45.07 7.30 41.53 23.79 19.81 11.36 6.45 3.11
D.4 Supplementary Ablation Study